The length of a rectangular room is two feet less than three times its width. The area of the
rectangle is 176 square feet.
a) Write an algebraic equation to model this problem.
b) Solve the equation to find the dimensions of the rectangle.

Answer by Boreal(15235)

x=length
y=width
xy=176
x=3y-2
(3y-2)*y=3y^2-2y=176
3y^2-2y-176=0
(3y+2)(y-8)=0
positive root is y=8 feet
x=22 feet, 3y-2
area is 176 ft^2.